An executive committee consists of 15 members, 7 men and 8 women. 4 members are selected at random to attend a meeting in Hawaii. The names are drawn from a hat. What is the probability that all 4 selected are women.
Thank you

Answer by ewatrrr(24785)

 
Hi,
committee of 15 members, 7 men and 8 women.
4 selected at random to attend a meeting in Hawaii
P(all women) = 8C4 / 15C4 = 70/1365 = .0513, 5.13% chance of doing that.